OpenStudy (cheerleader):

$13,543.48. $12,876.47. $14,064.21. $15,135.09. None of these choices are correct.

OpenStudy (cheerleader):

@pooja195

OpenStudy (sshayer):

\[FV=9000\left( 1+\frac{ 7.5 }{ 400 } \right)^{4 \times( 5.5)}=?\]

OpenStudy (sshayer):

\[=9000\left( \frac{ 4075 }{ 4000 } \right)^{22}=9000\left( \frac{ 163 }{ 160 } \right)^{22}=?\]

OpenStudy (sshayer):

use your calculator or log tables to find amount.

OpenStudy (mathmate):

@sshayer Not sure if OP even knows what a log table is! lol